OpenStudy (anonymous):

(a^5b^3) (a^4b^5) Do I multiply for add the exponents? E.g. would I have a^9 or a^20? Thank you!

jimthompson5910 (jim_thompson5910):

You add the exponents in this case

OpenStudy (anonymous):

Oh my goodness... THANK YOU!!!!! Last time I had to ask this question, I was forced to wait 10 minutes for a wrong answer.

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

So it is a^9andb^8 correct?

jimthompson5910 (jim_thompson5910):

bingo

OpenStudy (anonymous):

THANK YOU! Can you answer one more quick question, please?

jimthompson5910 (jim_thompson5910):

so \[\Large a^9b^8\]

OpenStudy (anonymous):

(-2hi^3)(2h^2ij^3)

jimthompson5910 (jim_thompson5910):

multiply -2 and 2 to get?

OpenStudy (anonymous):

Would it be -4h^3i^4j^7?

jimthompson5910 (jim_thompson5910):

The term for j is j^3 since the first expression doesn't have any j terms

jimthompson5910 (jim_thompson5910):

or because (-2hi^3)(2h^2ij^3) really is (-2hi^3j^0)(2h^2ij^3)

jimthompson5910 (jim_thompson5910):

So i'm not sure how you're getting 4+3 = 7

OpenStudy (anonymous):

-4h^2i^3j^12?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

What are the exponents for j in (-2hi^3j^0)(2h^2ij^3) ???

OpenStudy (anonymous):

Do I not multiply -2 and 2?

jimthompson5910 (jim_thompson5910):

yes those are the coefficients

OpenStudy (anonymous):

0 and 3

jimthompson5910 (jim_thompson5910):

add them

jimthompson5910 (jim_thompson5910):

to get the final exponent for j

OpenStudy (anonymous):

3

jimthompson5910 (jim_thompson5910):

So the answer is \[\Large -4h^3i^4j^3\]

OpenStudy (anonymous):

Thank you!!

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

Okay now, another one! :D ([3^2]^3g^5h^8)^2

jimthompson5910 (jim_thompson5910):

What is [3^2]^3

OpenStudy (anonymous):

Do you mind if I do the problem by myself, and I'll give you my answer and see if I'm right?

jimthompson5910 (jim_thompson5910):

alright

OpenStudy (anonymous):

Stay with me, please! :)

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

729g^10h^16

OpenStudy (anonymous):

Yes?

jimthompson5910 (jim_thompson5910):

no it's not correct

OpenStudy (anonymous):

Huh..

jimthompson5910 (jim_thompson5910):

3^2 is 9 So [3^2]^3 = 9^3 = 729 This means ([3^2]^3g^5h^8)^2 becomes (729g^5h^8)^2

jimthompson5910 (jim_thompson5910):

did you get that as one of your steps?

OpenStudy (anonymous):

Uhm

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

Then you square everything inside

OpenStudy (anonymous):

Oh!

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

531,411g^10h^13

OpenStudy (anonymous):

16

jimthompson5910 (jim_thompson5910):

yes, but I would get rid of the comma....computer answer systems don't like commas

jimthompson5910 (jim_thompson5910):

oh yes, 16 not 13

jimthompson5910 (jim_thompson5910):

use only commas to separate out answers (like ordered pairs), don't enter commas for large numbers

OpenStudy (anonymous):

kk

jimthompson5910 (jim_thompson5910):

so the answer is 531411g^10h^16 which looks like \[\Large 531411g^{10}h^{16}\]

OpenStudy (anonymous):

uhm

jimthompson5910 (jim_thompson5910):

what's wrong?

OpenStudy (anonymous):

x(x^4)(x^6)

jimthompson5910 (jim_thompson5910):

x = x^1

OpenStudy (anonymous):

x^11?

jimthompson5910 (jim_thompson5910):

So x(x^4)(x^6) is the same as x^1(x^4)(x^6) or x^1 times x^4 times x^6

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

x(x^4)(x^6) = x^11

OpenStudy (anonymous):

Okay, one more question! :)

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):